OpenStudy (anonymous):

F(x)=x to the 3rd power f(-2)=??

OpenStudy (anonymous):

-2^3 = -2*-2*-2 = -8

OpenStudy (anonymous):

just sub x=-3 into the function\[f(x)=x^3\]\[f(-2)=(-2)^3=-8\]

OpenStudy (anonymous):

i am confused how do you get 8?

OpenStudy (anonymous):

-8 i mean

OpenStudy (anonymous):

oh nvm -2x-2x-2

OpenStudy (anonymous):

raiderjoey showed you the definition of (-2)^3; it means 3 factors of -2, ie, (-2)(-2)(-2)=-8

OpenStudy (anonymous):

yup you got it

OpenStudy (anonymous):

thanks you guys :D i appreciate it

OpenStudy (anonymous):

k